The Surf-excel company uses a machine to fill boxes with soap powder. Assume that the net weight of the boxes of soap is normally distributed with mean 15 ounces and standard deviation 0.8 ounces. 25% of the boxes will be heavier than a certain net weight w and 75% of the boxes will be lighter than this net weight. Find w

1
Expert's answer
2022-02-24T17:50:37-0500

Let "W=" the net weight of the box: "W\sim N(\mu, \sigma)."

Given "\mu=15\ ounces, \sigma=0.8\ ounces."


"P(W<w)=P(Z<\dfrac{w-15}{0.8})=0.75"

"\dfrac{w-15}{0.8}\approx0.67449"

"w=15.54\ ounces"